A fish weighs 2LBS and it is 1/8 of what a chinchilla weighs. Find the equation and solve.
Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

2/(1/8)
2*8/1=16/1=16 lbs

Answer by richwmiller(17219)

Definitely something wrong here.
Chinchillas usually weigh about 1.3-4 pounds.
2=1/8c
c=16
no way!